Marginal tax rates   Partner​ A, a single​ taxpayer, is one of two partners in a small business. As​ such, she receives​ pass-through income that is taxed at her personal tax rates. After all adjustments and deductions have been​ made, including the​ 20% qualified business income​ deduction, she is preparing to calculate her taxes owed for the year. Using the tax rate schedule given here perform the​ following:

a. Find the marginal tax rate for the following levels of sole proprietorship earnings before​ taxes:

$15,400​

$60,600​

$91,500​

$146,000​

$255,000​

$452,000​

$1.4 million

b. Plot the marginal tax rates​ (measured on the y-​axis) against the pretax income levels​ (measured on the x-​axis). Explain the relationship between these variables.
Answer #1

Answer:

A) Marginal tax rate for sole proprietorship earnings before taxes are as follows:

A.1) Marginal tax rate for sole proprietorship earnings before taxes of $15,400 is 12%

Hence, total taxes due for earnings before taxes of $15,400 = $953+(12% X ($15,400-$9,525)) = $1,658

Note : 12% X amount over $9,525

A.2) Marginal tax rate for sole proprietorship earnings before taxes of $60,600 is 22%

Hence, total taxes due for earnings before taxes of $60,600 = $4,454+(22% X ($60,600-$38,700)) = $9,272

Note: 22% X amount over $38,700

A.3) Marginal tax rate for sole proprietorship earnings before taxes of $91,500 is 24%

Hence, total taxes due for earnings before taxes of $91,500 = $14,090+(24% X ($91,500-$82,500)) = $16,250

Note: 24% X amount over $82,500

A.4) Marginal tax rate for sole proprietorship earnings before taxes of $146,000 is 24%

Hence, total taxes due for earnings before taxes of $146,000 = $14,090+(24% X ($146,000-$82,500)) = $29,330

Note: 24% X amount over $82,500

A.5) Marginal tax rate for sole proprietorship earnings before taxes of $255,000​ is 35%

Hence, total taxes due for earnings before taxes of $255,000​ = $45,690+(35% X ($255,000​-$200,000)) = $64,940

Note: 35% X amount over $200,000

A.6) Marginal tax rate for sole proprietorship earnings before taxes of $452,000​ is 35%

Hence, total taxes due for earnings before taxes of $452,000​ = $45,690+(35% X ($452,000​​-$200,000)) = $133,890

Note: 35% X amount over $200,000

A.7) Marginal tax rate for sole proprietorship earnings before taxes of $1.4 million is 37%

Hence, total taxes due for earnings before taxes of $1.4 million = $150,690+(37% X ($1,400,000​​-$500,000)) = $483,690

Note: 37% X amount over $500,000

B) The chart is given below.

Marginal Tax Rate (%) VS Pretax Income Level Marginal Tax Rate (%) Pretax Income Level $9,525 $38,700 $82,500 $157,500 $200,0

From the above chart it can be concluded that as income increases tax rate also increases. However, over $500,000 it remains constant as 37%.